OpenStudy (anonymous):

A country's population in 1992 was 50 million. In 2002 it was 53 million. Estimate the population in 2010 using the exponential growth formula. Round your answer to the nearest million.

OpenStudy (anonymous):

you function will look like \[A=50e^{rt}\] and you know if \[t=8\] then \[A=53\] so write \[53=50e^{3r}\] and solve for r

OpenStudy (anonymous):

\[50e^{3r}=53\] \[e^{3r}=\frac{53}{50}\] \[3r=ln({\frac{53}{50}})\] \[r=\frac{ln(\frac{53}{50})}{3}\]

OpenStudy (amistre64):

50 * (e^((1 / (2002 - 1992)) * ln(53 / 50) * (2010 - 1992))) = 55.5290907

OpenStudy (anonymous):

CORRECT!